OpenStudy (anonymous):

7th grade math please help! 3 1/2 + w - 5 1/4 = 8 Explain how to solve this equation pls

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (nerdsarecool):

So to solve it

OpenStudy (nerdsarecool):

You have to go from right to left

OpenStudy (nerdsarecool):

So

OpenStudy (nerdsarecool):

31/2 + w

OpenStudy (nerdsarecool):

Then

OpenStudy (anonymous):

wait hold on how did you get 31/2

OpenStudy (nerdsarecool):

Oops

OpenStudy (nerdsarecool):

I meant

OpenStudy (nerdsarecool):

3 1/2

OpenStudy (nerdsarecool):

Sorry

OpenStudy (anonymous):

oh ok

OpenStudy (nerdsarecool):

Then you do the answer to that minus 5 1/4

OpenStudy (nerdsarecool):

Equals 8

OpenStudy (nerdsarecool):

Thats how you solve it

OpenStudy (anonymous):

Well um can you try to explain it a bit better?

OpenStudy (nerdsarecool):

Ok

OpenStudy (nerdsarecool):

First you do

OpenStudy (nerdsarecool):

3 1/2 + w

OpenStudy (nerdsarecool):

Which gives you an answer

OpenStudy (nerdsarecool):

Then you use that answer

OpenStudy (anonymous):

what do you get?

OpenStudy (nerdsarecool):

I don't know

OpenStudy (nerdsarecool):

3 1/2w

OpenStudy (nerdsarecool):

Thats what you get

OpenStudy (anonymous):

ok, then what?

OpenStudy (nerdsarecool):

Then you use that subtract it by 5 1/4

OpenStudy (sleepyjess):

Okay, first we'll need to subtract \(3 \dfrac12 - 5\dfrac14\) To do that, we'll need to find a common denominator. Do you know how to do that?

OpenStudy (nerdsarecool):

4

OpenStudy (nerdsarecool):

4

OpenStudy (sleepyjess):

Nerds, I'm trying to get the asker to the answer

OpenStudy (anonymous):

Eh well it's four...

OpenStudy (sleepyjess):

Good! So how do we make 2 go to 4?

OpenStudy (anonymous):

multiply by 2

OpenStudy (anonymous):

.

OpenStudy (sleepyjess):

Yes! So we have to multiply the top AND bottom by 2, which gives us \(3 \dfrac24 - 5\dfrac14\) Now, what is 3-5? \(\dfrac24 - \dfrac14\)?

OpenStudy (anonymous):

-2

OpenStudy (sleepyjess):

Yes, and \(\dfrac24 - \dfrac14\)?

OpenStudy (anonymous):

1/4

OpenStudy (sleepyjess):

Yes! So now we have \(-2 \dfrac14 +w = 8\). Now, we need to find w, do you know how to do that?

OpenStudy (anonymous):

subtract -2 1/4 by 8?

OpenStudy (sleepyjess):

Close, -2 1/4 is a negative, so we need to add 2 1/4 to both sides :)

OpenStudy (anonymous):

oh

OpenStudy (sleepyjess):

So, 2 1/4 - 2 1/4 cancels out, leaving us with w, the 8 + 2 1/4 =?

OpenStudy (anonymous):

72/4 ?

OpenStudy (sleepyjess):

Not quite, 8 + 2 = 10, 10 + 1/4 = 10 1/4

OpenStudy (sleepyjess):

So now we have, w = \(10 \dfrac14\)

OpenStudy (anonymous):

oh so that's it? the k12 answer thing says 9 3/4 :o

OpenStudy (sleepyjess):

that's weird...

OpenStudy (sleepyjess):

@TheSmartOne , did I do something wrong?

pooja195 (pooja195):

Yes you did

OpenStudy (sleepyjess):

Where?

pooja195 (pooja195):

Turn the fractions into improper fractions first

OpenStudy (sleepyjess):

7/2 21/4

OpenStudy (anonymous):

thats what I did and I got a totally different answer

OpenStudy (sleepyjess):

14/4 - 21/4 7/4 1 3/4

OpenStudy (sleepyjess):

Damn.... I'm sorry :(

OpenStudy (anonymous):

@sleepyjess We all make mistake :D But how do I get the answer?

pooja195 (pooja195):

Ok 1) Turn them into improper 7/2 21/4 \[\huge~\rm~ \frac{ 7 }{ 2}+w-\frac{ 21 }{ 4 } \]

pooja195 (pooja195):

\[\huge~\rm~\frac{ 7 }{ 2 }-\frac{ 21 }{ 4 }=\frac{ -7 }{ 4 }\]

OpenStudy (anonymous):

what happened to the addition sign?

pooja195 (pooja195):

The addition sign is still there however we can take out the w for the moment and we would subtract the 2 fractions

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I divided -7 and 4 and I got -1 3/4

pooja195 (pooja195):

\[\huge~\rm~\bf~w+\frac{- 7 }{ 4 }+\frac{ 7 }{ 4 }=8+\frac{ 7 }{ 4 }\] can you solve for w you dont divide those leave it as it is

OpenStudy (anonymous):

Ok so I got 15/4?? Is that wrong

pooja195 (pooja195):

The 7/4's on the w side cancel out 15/4 is incorrect So you really only need to do : \[\huge~\rm~\bf~8+\frac{ 7 }{ 4}=?\]

OpenStudy (anonymous):

Yes I did that and got 15/4

OpenStudy (anonymous):

wait I read that wrong

pooja195 (pooja195):

Its alright try again :-)

OpenStudy (anonymous):

Can you explain how to do that I keep getting the wrong answer

pooja195 (pooja195):

Ok First we need to find a common denominator

pooja195 (pooja195):

*least common

OpenStudy (anonymous):

4 right?

pooja195 (pooja195):

Yes 4 is right \[\huge~\rm~\bf~\frac{ (4)8 }{ (4)1 }+\frac{ 7 }{ 4 }\] \[\huge~\rm~\bf~\frac{ 32 }{ 4 }+\frac{ 7 }{ 4 }\] \[\huge~\rm~\bf~\frac{ 32+7=? }{ 4}\]

OpenStudy (anonymous):

39/4

OpenStudy (anonymous):

Oh got the answer thank you Pooja

pooja195 (pooja195):

you're welcome ^_^